1/(xy²+y⁴) is an integrating factor for the differential equation (x²y+y²)dx+(y³-x³)dy=0. Is it true or false. Give reasons for your answer


1
Expert's answer
2021-02-09T08:52:25-0500

"\\displaystyle\n\\textsf{It is not an integrating factor}\\\\\n(x^2 y + y^2)\\mathrm{d}x + (y^3 - x^3) \\mathrm{d}y = 0\\\\\n\n\\frac{\\mathrm{d}y}{\\mathrm{d}x} + \\frac{(x^2 + y)y}{y^3 - x^3} = 0\\\\\n\n\\frac{\\mathrm{d}}{\\mathrm{d}x}\\left(\\frac{y}{xy^2 + y^4}\\right) \\neq \\frac{\\mathrm{d}y}{\\mathrm{d}x} + \\frac{(x^2 + y)y}{y^3 - x^3} \\\\\n\n\n\\textsf{Or}\\\\\n(x^2 y + y^2)\\mathrm{d}x + (y^3 - x^3) \\mathrm{d}y = 0\\\\\ny(x^2\\mathrm{d}x + y^2\\mathrm{d}y) + y^2\\mathrm{d}x - x^3\\mathrm{d}y = 0\\\\\n\\textsf{Multiply both sides by}\\,\\, \\frac{1}{xy^2 + y^4} \\\\\n\n\\frac{x\\mathrm{d}x + y^2\\mathrm{d}y}{y(x + y^2)} + \\frac{y^2\\mathrm{d}x - x^3\\mathrm{d}y}{y^2(x + y^2)} = 0\\\\\n\n\\frac{\\mathrm{d}\\left(\\frac{x^2}{2} + \\frac{y^3}{3}\\right)}{y(x + y^2)} + \\frac{y^2\\mathrm{d}x - x^3\\mathrm{d}y}{y^2(x + y^2)} = 0\\\\\n\n\t\n\\textsf{This cannot be simplified further}\\\\\n\\textsf{to obtain a solution of the ODE}\\\\\n\n\\textsf{Therefore}\\,\\, \\frac{1}{xy^2 + y^4}\\,\\,\\textsf{is not an integrating factor}"